OpenStudy (anonymous):

INTEGRATE : 5x/[(4 + x^2)]^2

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 5x }{ (4+x ^{2})^{2} }\]

OpenStudy (anonymous):

(5 x)/(x^2+4)^2

OpenStudy (anonymous):

Did that help?

OpenStudy (amistre64):

what is the derivative of the bottom?

OpenStudy (amistre64):

16x + 4x^2 seems about right for the derivative 5x + (11x+4x^2) - (11x+4x^2) doesnt seem to work out well might try a different method maybe some usub?

OpenStudy (amistre64):

u = 4+x^2 du = 2x dx ... du/2 = x dx

OpenStudy (amistre64):

so\[\int\frac{5}{u^2}*\frac{du}{2}\]

OpenStudy (anonymous):

sorry guys this thing went "updating till now! I am trying to follow your work

OpenStudy (anonymous):

and thank you @amistre64, the answer is very right!